OpenStudy (erinweeks):

2 more questions tonight!

OpenStudy (erinweeks):

8 + 16 + 24 + . . . + 8n = 4n(n + 1)

OpenStudy (shamil98):

let it is true for n=1 so 8(1)=4(1)(1+1) 8=8 fist condition is true now let its true for n=k 8 + 16 + 24 + . . . + 8k = 4k(k + 1) we have to show that it is true for n=k+1

OpenStudy (erinweeks):

okay how we do that

OpenStudy (shamil98):

when n=k+1 we need to show that 8 + 16 + 24 + . . . + 8k +8(K+1)=4(k+1)((k+1)+1) 8 + 16 + 24 + . . . + 8k +8(K+1)= 4k(k + 1)+8(K+1).....1 because 8 + 16 + 24 + . . . + 8k=4k(k+1) simplify the 1 =4k(k + 1)+8(K+1) =4k(k+1)+8k+8 =4k^2+4k+8k+8 =4k^2+12k+8 take 4 common =4(k^2+3k+2) =4(k+1)((k+2) or =4(k+1)((k+1)+1) Your question i believe is using mathematical induction to prove that that statement is positive for every positive integer n.

OpenStudy (erinweeks):

im lossst

OpenStudy (shamil98):

@wio takeover please. I'm a bit tired D:

OpenStudy (erinweeks):

@ganeshie8 please help me you usuallyy make me undertsand

OpenStudy (shamil98):

That proves your statement btw.